OpenStudy (danielbarriosr1):

Which equation represents the line passing through the points (3,2) and (-9,6)

OpenStudy (danielbarriosr1):

Possible answers: A. x-3y=9 B. x+3y=9 C. 3x-y=-9 D. 3x+y=9

OpenStudy (danielbarriosr1):

@mathmale

OpenStudy (mathmale):

Hola, Daniel! Que va? Como estas?

OpenStudy (danielbarriosr1):

Bien y tu?

OpenStudy (mathmale):

I would take those two points, find the slope (m), and then write the equation of the line in point-slope form. Bien! Gracias.

OpenStudy (danielbarriosr1):

m=0.33333

OpenStudy (mathmale):

Bet you already know the formula for slope:\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (danielbarriosr1):

sorry -0.33333

OpenStudy (mathmale):

and also that you know the point-slope form:\[y-y _{0}=m(x-x _{0})\]

OpenStudy (mathmale):

Daniel: Think: As you go from (-9,6) to (3,2), it does seem that y is decreasing as x increases. Therefore, your negative slope makes sense!

OpenStudy (danielbarriosr1):

yes

OpenStudy (mathmale):

You may use either -0.333333 or -1/3 for the slope, m. I'd use m=-1/3 myself. Please choose either point (it doesn't matter which one). Substitute the coordinates of that chosen point into the point-slope form of the eq'n of a straight line, along with your value of m. What do you get?

OpenStudy (mathmale):

Once you have this result, you need only re-arrange the terms so that x comes first and y second in the "equation of a line in 'standard form' ", with a constant on the right.

OpenStudy (acxbox22):

slope is -1/3 select either point and substitute the points x and y values into the point slope form equation just like @mathmale said i would choose (3,2) as it is smaller y-2=-1/3(x-3) now just change it into slope intercept form

OpenStudy (danielbarriosr1):

y=-0.3x+2.9

OpenStudy (mathmale):

Daniel, please multiply both sides of your result by 10 to eliminate the decimal fractions. Then rearrange your equation so that you have 3x on the left, followed by 1y, followed by an equal sign and a constant.

OpenStudy (acxbox22):

nvr mind sorry i forgot to look at the answer choices change it to standard form and use -1/3 for slope

OpenStudy (danielbarriosr1):

10y=-3x+29

OpenStudy (danielbarriosr1):

10y+3x=29

OpenStudy (mathmale):

or 3x + 10y = 29. Looks very nice. Does this match any of the four possible answers?

OpenStudy (danielbarriosr1):

nop

OpenStudy (mathmale):

Darn. I'm going to go back and review what you're done in the hope of being able to point out what went wrong, if anything. In the meantime, why not look at the next homework problem you need to do?

OpenStudy (danielbarriosr1):

OK

OpenStudy (anonymous):

just check on the graphing calculator https://www.desmos.com/calculator

OpenStudy (acxbox22):

y-2=-1/3(x-3) y-2=-1/3x+1 y=-1/3x+3 1/3x+y=3 x+3y=9

OpenStudy (acxbox22):

the answer is b

OpenStudy (anonymous):

it is b i checked on the graph website

OpenStudy (texaschic101):

you didn't have to use point slope form

OpenStudy (mathmale):

Confirming that your slope, m, is correct: m = -1/3. Next time, share what you've done towards finding the equation of the line. I would choose the point (3,2). Then x=3 and y=2. Substituting these, and m = -1/3, into the point-slope formula results in \[y-2=\frac{ -1 }{ 3 }(x-3)\]Daniel: Would you plese simplify this? the first thing I'd do would be to multiply everything in this equation by 3, to remove the fractions.

OpenStudy (acxbox22):

i already simplified it @danielbarriosr1

OpenStudy (texaschic101):

you could have just used y = mx + b form

OpenStudy (mathmale):

There are, of course, various ways in which to do this problem; I tend to use the point-slope form because it's so easy and straightforward. It's also easy to change from one form to another that way. Daniel? How are you doing in terms of simplifying this equation?

OpenStudy (danielbarriosr1):

Im on it

OpenStudy (mathmale):

Several of you are trying to help at the same time. If you have an alternative approach, and see nothing wrong with the one I'm using, I ask that you please wait until Daniel is finished BEFORE proposing an alternative approach. Also, remember, the goal is for Daniel to learn how to do these problems himself, not for you to provide answers. Thanks.

OpenStudy (danielbarriosr1):

y=-1/3x+3

OpenStudy (mathmale):

Daniel, here's what I was suggesting earlier: Please multiply the following equation (that is, each term of the following equation) by 3 to eliminate the fraction:\[3[y-2=\frac{ -1 }{ 3 }(x-3)\]

OpenStudy (mathmale):

The result is 36 - 6 = -(x-3). Are you comfortable with this, Daniel?

OpenStudy (danielbarriosr1):

3y-6=-1(x-3)

OpenStudy (danielbarriosr1):

where does the 36 came from?

OpenStudy (acxbox22):

yeah?

OpenStudy (mathmale):

Typo. Your version is correct. Proceed, please.

OpenStudy (danielbarriosr1):

3y-6=-x+3

OpenStudy (danielbarriosr1):

3y=-x+9

OpenStudy (mathmale):

Great. Please move that x to the left side of your equation, and then youi'll be done.

OpenStudy (danielbarriosr1):

y+1/3x=3

OpenStudy (mathmale):

You had 3y=-x+9. I'd suggest you'd find it easier if you'd please just add x to both sides. NO division. Why: Because there are NO fractions in any of the four possible answers.

OpenStudy (danielbarriosr1):

oh

OpenStudy (mathmale):

So that results in x+3y=?

OpenStudy (mathmale):

Is your result the same as any of the four possible answers?

OpenStudy (danielbarriosr1):

x+3y=9

OpenStudy (danielbarriosr1):

Yes... B.

OpenStudy (danielbarriosr1):

Thank you @mathmale

OpenStudy (mathmale):

Cool! Congrats!! @acxbox22 claims that he has a faster approach. Daniel, it's entirely up to you to choose whether you want him/her to present that approach or to let you move on to the next homework problem. De nada!

OpenStudy (danielbarriosr1):

and thank you @acxbox22 of course

OpenStudy (acxbox22):

it is allright as i already showed him my approach 15 posts earlier

OpenStudy (danielbarriosr1):

But I prefer @mathmale 's explanations

OpenStudy (acxbox22):

no problem

OpenStudy (mathmale):

Thank you, Dan. I'd like to point out that acxbox22 did obtain the correct result with less work, but essentially he followed the same procedure as you and I did. I just prefer to explain problem solutions verbally instead of with just equations. Suggest we move on immediately to your next homework problem, if there is one. @acxbox22: Sincere thanks for your contributions here.

OpenStudy (danielbarriosr1):

Ok...

OpenStudy (danielbarriosr1):

Which equation represents the line passing through the point (4,-16) and is perpendicular to the line...

OpenStudy (danielbarriosr1):

perpendicular to the line y=-2/3x+8

OpenStudy (danielbarriosr1):

possible answers: A. y=2/3x-22 B. y=-2/3x-22 C. y=-3/2x-22 D. y =3/2x-22

OpenStudy (acxbox22):

do you know perpendicular lines?

OpenStudy (danielbarriosr1):

I know I can eliminate B and C because that slope would make them parallel

OpenStudy (danielbarriosr1):

Those are the lines that intersect at one point

OpenStudy (danielbarriosr1):

I think is D.

OpenStudy (acxbox22):

you can eliminate B but not C since it does not have the same slope in fact you can find out the answer without much work

OpenStudy (danielbarriosr1):

But C is negative

OpenStudy (acxbox22):

perpendicular means opposite reciprocal what is the opposite reciprocal of -2/3? if you find that out then you will know the answer since each answer choice's slopes are different

OpenStudy (danielbarriosr1):

3/2

OpenStudy (danielbarriosr1):

D right?

OpenStudy (acxbox22):

exactly.. d is your answer

OpenStudy (danielbarriosr1):

Thanks

OpenStudy (acxbox22):

np

OpenStudy (anonymous):

it is b

OpenStudy (acxbox22):

no b would make it parallel not perpendicular

OpenStudy (danielbarriosr1):

It is supose to be perpendicular no parllel

OpenStudy (anonymous):

oh